OpenStudy (anonymous):

Patrycia’s overall grade in her physics class is based solely on the average of her scores on 6 unit tests. The grade she received on the first 5 tests are as follows: 89, 93, 95, 92, 90. What grade does Patrycia need to earn on her 6th unit test to end up with a 90% in the class?

OpenStudy (texaschic101):

(89 + 93 + 95 + 92 + 90 + x)/6 = 90% (459 + x)/6 = 90% (multiply by 6) 459 + x = 540 x = 540 - 459 x = 81 check... (459 + 81)/6 = 90% 540/6 = 90% 90% = 90% (correct) She would have to make an 81